OpenStudy (anonymous):

I chose B and C. is that correct? 2. Determine which ordered pair in the form (a, b) is a solution of the given equation. 7a – 5b = 28 (Points : 1) (–3, –2) (–2, –3) (0, 4) (4, 0) 3. Determine whether the ordered pairs in the form (x, y) are solutions of the given equation and select the correct answer. (Points : 1) Yes, both are solutions. No, neither is a solution. The first is a solution, but the second is not. The first is not a solution, but the second is.

OpenStudy (anonymous):

I CHOSE B & D is that correct 4. Determine which ordered pair in the form (x, y) is a solution of the given equation. (x + 7)y = 10 (Points : 1) (2, 1) (3, 1) (4, 1) (13, 2) 5. Which pairs are solutions to the equation? 3xy + 9 = 33 (Points : 1) (1, 8) and (8, 1) (1, 8) and (1, 11) (1, 11) and (2, 6) (8, 1) and (2, 6)

OpenStudy (anonymous):

Number two b no that is completely wrong

OpenStudy (anonymous):

You want the answer

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

It is D for the one labeled 2

OpenStudy (anonymous):

ok and the rest are correct?

OpenStudy (anonymous):

7*4-5*0=28

OpenStudy (anonymous):

um let me check

OpenStudy (anonymous):

On number 3 i dont see an equation

OpenStudy (anonymous):

On number 4 b is correct

OpenStudy (anonymous):

6x-y/2=14 (-9,3), (2,-4)

OpenStudy (anonymous):

And that is correct

OpenStudy (anonymous):

Anything else you need help on

OpenStudy (anonymous):

is 5 correct?

OpenStudy (anonymous):

What grade math is this

OpenStudy (anonymous):

Let me check

OpenStudy (anonymous):

10th

OpenStudy (anonymous):

No 5 is a but im in 8th and these seem easy to me and i usually suck at math but i just got these problems i don't know why.

OpenStudy (anonymous):

If you need help with anything else send me amessage

OpenStudy (anonymous):

i suck at algebra but i am taking geometry its way easier

OpenStudy (anonymous):

Oh thats ok

OpenStudy (anonymous):

you're smart thanks for your help! :)

OpenStudy (anonymous):

Thank you but did i just send a message twice

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Oh ok thanks

OpenStudy (anonymous):

@FateRocks8D Helpp !